For the following function, determine its inverse f-1. Use the switch and solve strategy:
f(x) = 3x - 8

Answer by eperette(173)

f(x) = 3x - 8 or y=3x-8
(1)switch x with y
x = 3y - 8
(2) solve for y
x = 3y - 8
x + 8 = 3y - 8 + 8
x + 8 = 3y + 0
(x + 8)/3 = 3y/3
(x + 8)/3 = y
or
f^-1(x)= (x + 8)/3